OpenStudy (anonymous):

Four students share $100. Al and Bob get 1/10 and 1/5 of the total respectively. Carl gets the average of what Al and Bob got. Dave gets the rest. how much, in dollars , did Dave get?

OpenStudy (anonymous):

first, you know Dave gets the rest, that's [1-(what the other three get)] times 100

OpenStudy (anonymous):

we have what Al and Bob get. what Carl gets is their average which is: \[{1/10 + 1/5 \over 2} = {3/10 \over 2}=3/20\]

OpenStudy (anonymous):

just add what the three get: 3/20+1/10+1/5=0.45. finally, Dave share is 1-0.45=0.55.. in dollars, he gets 0.55*100= $55

OpenStudy (anonymous):

there is an easier way to solve it, if you think this way it a bit complicated.

OpenStudy (anonymous):

thank you :) !

OpenStudy (anonymous):

you're welcome :)